OpenStudy (anonymous):

During the past school year, there were 640 students and 52 teachers. What is the ratio of students to each teacher rounded to the nearest tenth?

OpenStudy (anonymous):

640/52=12.307... 12.3:1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Cross-multiplication is helpful in

OpenStudy (anonymous):

There are 96 girls in a junior class of 176 students. Find the ratio of girls to boys. a. 11:6 c. 6:11 b. 5:6 d. 6:5

OpenStudy (anonymous):

c

OpenStudy (anonymous):

All you would do is subtract the girls from 176, so 176-96=80 since the ratio is G:B its 96:80 96/8=12 80/8=10 simplify: the ratio is 6:5

OpenStudy (anonymous):

Its C